OpenStudy (anonymous):

How would you find the exact value of cos(sin^-1(5/13) -cos^-1(4/5))? Does this require a trig identity?

OpenStudy (anonymous):

For similar problems, I had drawn a triangle. But these don't have the same hypotenuses, so I am confused.

OpenStudy (zarkon):

\[\cos(\alpha-\beta) = \cos(\alpha)\cos(\beta)+\sin(\alpha)\sin(\beta)\]

OpenStudy (zarkon):

then do two triangles

OpenStudy (anonymous):

is the answer 63/65

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

Good thank you